OpenStudy (anonymous):

find the derivative of the function f(x) = x(3x-12)^4

OpenStudy (anonymous):

product plus chain rule

OpenStudy (anonymous):

Product rule: (derivative of (x)) * (3x-12)^4 * (derivative of (3x-12)^4)*x Use chain rule to get the derivative of (3x-12)^4

OpenStudy (anonymous):

See video here http://www.khanacademy.org/math/calculus/v/the-chain-rule

OpenStudy (inkyvoyd):

remember the chin rule: f(g(x))=f'(g(x)g'(x)

OpenStudy (inkyvoyd):

sometimes written as f(u)=f'(x)u'

OpenStudy (inkyvoyd):

(3x-12)^4 ->u=3x-12 f(u)=u^4

OpenStudy (anonymous):

find the derivative of the function f(x)=x(3x-12)^4 .. so f'(x) =?

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

See the video link I gave above. It is very helpful indeed